# An extension of Priestley duality to fuzzy topologies and positive MV-algebras

### v1

@@ Abstract
We extend Priestley Duality to suitable categories of fuzzy topological spaces and ordered algebraic structures that generalize bounded distributive lattices. The duality we prove extends not only classical Priestley Duality between Priestley Spaces and bounded distributive lattices, but also the duality between limit cut complete MV-algebras and Stone MV-topological spaces [15] which, on its turn, is an extension of classical Stone Duality.
@@ 2 Preliminaries
An MV-algebra is a structure $(A,\oplus,*,\mathbf{0})$ where $\oplus$ is a binary operation, * is a unary operation and $\mathbf{0}$ is a constant such that the following axioms are satisfied for any $a,b\in A$:
MV1)
$(A,\oplus,^{*},\mathbf{0})$ is an Abelian monoid,
MV2)
@@
We denote an MV-algebra $(A,\oplus,^{*},\mathbf{0})$ simply by its universe A, and define the constant $\mathbf{1}$ and the derived operations $\odot$ and $\ominus$ as follows:
  • $\mathbf{1}:=\mathbf{0}^{*}$
@@
Let $A$ an MV-algebra and $S$ a nonempty subset of $A$. Set $S_{0}=S$ and, for all $i\geq 1$, $S_{i}=S_{i-1}\cup S_{i-1}^{*}\cup\{a\oplus b:a,b\in S_{i-1}\}$. Then $\langle S\rangle=\bigcup\limits_{i\geq 0}S_{i}$.
@@
An ideal of an MV-algebra $A$ is a subset $I$, is downward closed submonoid of $(A\oplus,0)$. It is well-known that ideals and congruences of every MV-algebra are in bijective correspondence. In fact, for any congruence, the class of 0 is an ideal and, for every ideal $I$, the relation defined by $a\sim_{I}b$ if and only if $d(a,b):=(a\ominus b)\oplus(b\ominus a)\in I$ is a congruence and is the unique one whose 0-class is $I$.
@@
    simple if its only proper ideal is $\{0\}$
@@
[2] Every semisimple MV-algebra can be embedded in the MV-algebra of fuzzy subsets $0,1]^{\operatorname{Max}A}$ of $\operatorname{Max}A$.
@@
(Sketch) For each maximal ideal $M$ of $A$, the quotient algebra $A/M$ is simple and, therefore, isomorphic to a subalgebra of $[0,1]$. So, there exists an embedding $\iota_{M}:A/M\to[0,1]$. Let $\iota:A\to[0,1]^{\operatorname{Max}(A)}$ be the map defined by $\iota(a)=\hat{a}$ for each $a\in A$, where
$\hat{a}:\operatorname{Max}(A)$ $\to[0,1]$
@@
$\ker\iota=$ $\bigcap_{M\in\operatorname{Max}(A)}\{a\in A:\hat{a}(M)=0\}$
$=$ $\bigcap_{M\in\operatorname{Max}(A)}\{a\in A:\iota_{M}(a)=(a/M)=0\}$
$=$ $\bigcap_{M\in\operatorname{Max}(A)}M=\operatorname{Rad}(A).$
@@
[15, Proposition 4.5] Let $A$ be a semisimple MV-algebra. Then a cut $X$ of $A$ is a limit cut if and only if there exists a cut $Y$ of $A$ such that, in $0,1]^{\operatorname{Max}A}$, $\bigvee\widehat{X}=\bigwedge\widehat{Y}^{*}$, where $Y^{*}=\{y^{*}\mid y\in Y\}$. Moreover, $Y$ is a limit cut too.
@@
[15, Corollary 4.6] A semisimple MV-algebra $A$ is lcc if and only if, for all $X,Y\subseteq A$ and $\alpha\in[0,1]^{\operatorname{Max}A}$, $\alpha=\bigvee\widehat{X}=\bigwedge\widehat{Y}$ implies $\alpha\in\widehat{A}$.
@@
  • closed if $f^{\to}(\beta)\in\tau^{*}_{Y}$ for all $\beta\in\tau^{*}_{X}$
@@
Let us consider the MV-topology $[0,1]$ on a singleton $\{x\}$. For any $n>1$, all the sets of type $[0,1/n]$, $0,1/n]\cap\mathbb{Q}$, and $0,1/n]\setminus\mathbb{Q}$ are easily seen to be (non-large) subbases for the given MV-topology. Also $0,1]\setminus\mathbb{Q}$ is a subbase — it is a base, in fact.
@@
[15, Lemma 3.9] If $(X,\tau)$ is an Hausdorff space, then all crisp singletons of $X$ are closed.
@@ 3 MV-lattices and Positive MV-algebras
    $(x\cdot y)\vee(x\wedge y)=x\wedge y$
  4. 4.
    $(x+y)\vee(x\vee y)=x+y$
@@
. Every bounded distributive lattice, with $\vee$ and $\cdot:=\wedge$ , is an MV-lattice.
Given an MV-algebra $(A,\oplus,^{*},0)$ and considering its derived operations, the algebra $(A,\oplus,\odot,\vee,\wedge,0,1)$ is an MV-lattice. Any subalgebra $(B,\oplus,\odot,\vee,\wedge,0,1)$ of such an MV-lattice is called a positive subreduct of the MV-algebra $A$. The following definition is a reformulation of [1, Definition 2.2].
@@
In the case of MV-algebras $\text{\L }_{n}$, it was shown in [13] that all of their positive MV-subreducts are exactly the MV-subalgebras, i. e., the $\text{\L }_{k}$’s, for all divisor $k$ of $n$.
@@
The set $A=\left\{(0,0),\left(0,\frac{1}{2}\right),(0,1),\left(\frac{1}{2},1\right),(1,1)\right\}\subset\text{\L }_{2}\times\text{\L }_{2}$ is a generating positive MV-algebra of $\text{\L }_{2}\times\text{\L }_{2}$
@@
where, the maximal ideals of $\text{\L }_{2}\times\text{\L }_{2}$ are $I_{1}=\{(0,0),(0,\frac{1}{2}),(0,1)\}$ and $I_{2}=\{(0,0),(\frac{1}{2},0),$ (1,0)}(1,0)\}. Then, $\operatorname{Max}_{\leq}A=\{I_{1}^{A},I_{2}^{A}\}$ where $I_{2}^{A}=\{(0,0)\}\subset I_{1}=I_{1}^{A}$. So, $I_{2}^{A}$ is not a maximal ideal of $A$.
@@
The category of MV-lattices and its morphisms will be denoted by $^{\operatorname{MV}}\!\mathcal{L}$, and we will denote by $\mathcal{MV}_{+}$ and $\mathcal{MV}^{\operatorname{lcc}}_{+}$ its full subcategories of positive MV-algebras and $\operatorname{lcc}$ positive MV-algebras, respectively. Given an MV-lattice $(A,\vee,+,\cdot,\wedge,0,1)$, we consider the set of morphisms from $A$ to $[0,1]$, which is denoted by $H_{A}$ and is equipped with a partial order defined by $f\leq g$ if and only if $f(a)\leq g(a)$ for every $a$ in $A$. For $A$ as in Example 3.7, we have that $H_{A}=\{f_{1},f_{2}\}$ where $f_{i}:A\longrightarrow\text{\L }_{2}\subseteq[0,1]$ is defined by $f_{i}(x_{1},x_{2})=x_{i}$ for $(x_{1},x_{2})\in A$ and $f_{1}\leq f_{2}$. Note that $f_{1}$ and $f_{2}$ are restrictions of homomorphisms from the MV-algebra $\text{\L }_{2}\times\text{\L }_{2}$ to $[0,1]$. This is actually an instance of a general situation, as shown in the following lemma.
@@
The following result provides a necessary and sufficient condition for an MV-lattice $A$ to be embedded in a semisimple MV-algebra.
@@
It is known that an MV-algebra can be generated by various positive algebras. According to Proposition 3.9, the corresponding sets of homomorphisms from these algebras into $[0,1]$ are equipotent. However, these sets are not necessarily isomorphic partially ordered sets, as the following example shows.
@@
Given the MV-algebra $E=\text{\L }_{2}\times\text{\L }_{2}\times\text{\L }_{3}$, we have that the set
@@
    The positive MV-algebra
    $\begin{array}[]{ll}A=&\left\{\left(0,0,0\right),\left(0,\frac{1}{2},0\right),\left(0,1,0\right),\left(\frac{1}{2},1,0\right),\left(1,1,0\right),\right.\\ &\left.\left(0,\frac{1}{2},\frac{1}{3}\right),\left(0,1,\frac{1}{3}\right),\left(\frac{1}{2},1,\frac{1}{3}\right),\left(1,1,\frac{1}{3}\right),\left(0,1,\frac{2}{3}\right),\right.\\ &\left.\left(\frac{1}{2},1,\frac{2}{3}\right),\left(1,1,\frac{2}{3}\right),\left(0,1,1\right),\left(\frac{1}{2},1,1\right),\left(1,1,1\right)\right\}\end{array}$
@@
    The positive MV-algebra
    $C=B\cup\left\{\left(0,0,\frac{1}{3}\right),\left(0,0,\frac{2}{3}\right),\left(\frac{1}{2},\frac{1}{2},\frac{2}{3}\right),\left(0,\frac{1}{2},\frac{2}{3}\right),\left(0,0,1\right),\left(0,\frac{1}{2},1\right),\left(\frac{1}{2},\frac{1}{2},1\right)\right\}$
    is such that
    $\begin{array}[]{l}I_{1}^{C}=I_{1},\\ I_{2}^{C}=\left\{\left(0,0,0\right),\left(0,0,\frac{1}{3}\right),\left(0,0,\frac{2}{3}\right),(0,0,1)\right\},\text{ and}\\ I_{3}^{C}=\left\{(0,0,0),\left(0,\frac{1}{2},0\right),(0,1,0),\left(\frac{1}{2},\frac{1}{2},0\right),\left(\frac{1}{2},1,0\right),(1,1,0)\right\}.\end{array}$
    Moreover $f_{1_{C}}\leq f_{2_{C}}$ but $f_{2_{C}}$ and $f_{3_{C}}$ are not comparable, since
    $f_{3}\left(\frac{1}{2},\frac{1}{2},\frac{2}{3}\right)=\frac{2}{3}>\frac{1}{2}=f_{2}\left(\frac{1}{2},\frac{1}{2},\frac{2}{3}\right)\text{ and }f_{3}\left(\frac{1}{2},\frac{1}{2},0\right)=0<\frac{1}{2}=f_{2}\left(\frac{1}{2},\frac{1}{2},0\right).$
    $H_{C}$$\text{Max}_{\leq}(C)$$f_{2_{C}}$$I_{1}^{C}$$f_{3_{C}}$$I_{3}^{C}$$f_{1_{C}}$$I_{2}^{C}$}
    So, $H_{A}$ and $H_{C}$ are equipotent sets but are not order isomorphic.
@@
    If $x<y<\frac{1}{2}$, we apply Step 1 and we obtain that $y_{1}\ominus x_{1}\geq 2(y\ominus x)>y\ominus x$ where $y_{1}=ky$ and $x_{1}=kx$. Moreover, since $y_{1}>\frac{1}{2}$, then either $x_{1}\leq\frac{1}{2}$ or $x_{1}>\frac{1}{2}$. If $x_{1}\leq\frac{1}{2}$, by case 1, there is $n_{1}\in\mathbb{N}$ such that $n_{1}x_{1}^{2}=0$ and $n_{1}y_{1}^{2}=1$. That is, there is a term $t(a)=n_{1}(ka)^{2}$ such that $t(x)=0$ and $t(y)=1$. If $x_{1}>\frac{1}{2}$, we apply Step 2 and we obtain
    $y_{2}\ominus x_{2}=y_{1}^{h}-x_{1}^{h}=h(y_{1}\ominus x_{1})\geq 2(y_{1}\ominus x_{1})>y_{1}\ominus x_{1}$
    where $y_{2}=y_{1}^{h}$ and $x_{2}=x_{1}^{h}$ . Note that $x_{2}<\frac{1}{2}$ and again we can apply Step 1 or case 2 according to the possible values of $y_{2}$. Thus we define by recursion a sequence of pairs $(x_{i},y_{i})$ such that for each $i\geq 1$,  $y_{i+1}\ominus x_{i+1}\geq 2(y_{i}\ominus x_{i})>y_{i}\ominus x_{i}$. For this reason, for some $n\in\mathbb{N}$, $x_{n}$ and $y_{n}$ will satisfy the conditions of case 1 or 2 and we shall obtain a term $t$ such that $t(x)=0$ and $t(y)=1$.
@@ 4 The extension of Priestley Duality
    $x\nleq y$ implies that there is an increasing clopen set $\alpha$ such that $\alpha(x)=1$ and $\alpha(y)=0$.
@@
    $(X,\tau)$ has a basis of clopens
MV-spaces satisfying P2 are called totally order disconnected. It is easy to verify that every totally order disconected MV-space is a Hausdorff MV-space. We will denote by $^{\operatorname{MV}}\!\ \mathcal{T}\!\!\operatorname{op}_{\leq}$ the category of partially ordered MV-spaces whose morphisms are functions that are simultaneously continuous and order-preserving. Moreover, we shall denote by $^{\operatorname{MV}}\!\mathcal{P}\!\!\operatorname{ries}$ the full subcategory of Priestley MV-spaces.
@@
Let $(A,\vee,+,\cdot,\wedge,0,1)$ be an MV-lattice and $H_{A}$ be the set of homomorphisms $f:A\longrightarrow[0,1]$. $H_{A}$ has a partial order defined by $f\leq g$ iff $f(a)\leq g(a)$ for all $a\in A$. Now, in order to define an MV-topology on $H_{A}$, let us consider the set $\tilde{A}=\{\tilde{a}\mid a\in A\}$ as in 3.10, where
@@
The set $\tilde{A}$ generates a semisimple MV-algebra $B\subseteq[0,1]^{H_{A}}$ which forms a basis for an MV-topology $\tau_{A}$ on $H_{A}$. This partially ordered MV-space $(H_{A},\tau_{A},\leq)$ generated by $B$ is called the MV-space associated to $A$.
@@
${\operatorname{Clop}_{\uparrow}X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varepsilon_{\operatorname{Clop}_{\uparrow}X}}$$\scriptstyle{\operatorname{id}_{\operatorname{Clop}_{\uparrow}X}}$${\operatorname{Clop}_{\uparrow}X}$${\operatorname{Clop}_{\uparrow}\Upsilon\operatorname{Clop}_{\uparrow}X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Clop}_{\uparrow}\eta_{X}}$
${\Upsilon A\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta_{\Upsilon A}}$$\scriptstyle{\operatorname{id}_{\Upsilon A}}$${\Upsilon\operatorname{Clop}_{\uparrow}\Upsilon A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Upsilon\varepsilon_{A}}$${\Upsilon A}$
@@
For any MV-lattice $(A,\vee,\wedge,+,\cdot,0,1)$, the associated MV-space $(H_{A},\tau,\leq)$ is compact.
@@
Let $\mathcal{MV}^{\operatorname{lcc}}_{H}$ be the category of $\operatorname{lcc}$ positive MV-algebras that are $H$-complete and $^{\operatorname{MV}}\!\mathcal{P}\!\!\operatorname{ries}^{\operatorname{lcc}}$ be the category of Priestley MV-spaces whose set of increasing clopen sets is an $\operatorname{lcc}$ positive MV-algebra. Then, thanks to Theorem 4.5 and the last two lemmas, the following result readily follows.

### v2

@@ Abstract
We extend Priestley Duality to suitable categories of fuzzy topological spaces and positive MV-algebras, which stand to MV-algebras as bounded distributive lattices stand to Boolean algebras. The duality we prove extends not only classical Priestley Duality between Priestley Spaces and bounded distributive lattices, but also the duality between limit cut complete MV-algebras and Stone MV-topological spaces [14] which, on its turn, is an extension of classical Stone Duality.
@@ 2 Preliminaries
An MV-algebra is a structure $(A,\oplus,^{*},0)$ where $\oplus$ is a binary operation, is a unary operation and 0 is a constant such that the following axioms are satisfied for any $a,b\in A$:
$(A,\oplus,0)$ is an Abelian monoid,
@@
We denote an MV-algebra $(A,\oplus,^{*},0)$ simply by its universe A, and define the constant $1$ and the derived operations $\odot$ and $\ominus$ as follows:
  • $1:=0^{*}$
@@
Let $A$ be an MV-algebra and $S$ a nonempty subset of $A$. Set $S_{0}=S$ and, for all $i\geq 1$, $S_{i}=S_{i-1}\cup S_{i-1}^{*}\cup\{a\oplus b:a,b\in S_{i-1}\}$. Then $\langle S\rangle=\bigcup\limits_{i\geq 0}S_{i}$.
@@
An ideal of an MV-algebra $A$ is a subset $I$, is downward closed submonoid of $(A,\oplus,0)$. It is well-known that ideals and congruences of every MV-algebra are in bijective correspondence. In fact, for any congruence, the class of 0 is an ideal and, for every ideal $I$, the relation defined by $a\sim_{I}b$ if and only if $d(a,b):=(a\ominus b)\oplus(b\ominus a)\in I$ is a congruence and is the unique one whose 0-class is $I$.
@@
    simple if its only proper ideal is $\{0\}$, and
@@
[2] Every semisimple MV-algebra $A$ can be embedded in the MV-algebra of fuzzy subsets $0,1]^{\operatorname{Max}A}$ of $\operatorname{Max}A$.
@@
(Sketch) For each maximal ideal $M$ of $A$, the quotient algebra $A/M$ is simple and, therefore, isomorphic to a subalgebra of $[0,1]$. So, there exists an embedding $\iota_{M}:A/M\to[0,1]$. Let $\iota:A\to[0,1]^{\operatorname{Max}A}$ be the map defined by $\iota(a)=\hat{a}$ for each $a\in A$, where
$\hat{a}:\operatorname{Max}A$ $\to[0,1]$
@@
$\ker\iota=$ $\bigcap_{M\in\operatorname{Max}A}\{a\in A:\hat{a}(M)=0\}$
$=$ $\bigcap_{M\in\operatorname{Max}A}\{a\in A:\iota_{M}(a)=(a/M)=0\}$
$=$ $\bigcap_{M\in\operatorname{Max}A}M=\operatorname{Rad}A.$
@@
[14, Proposition 4.5] Let $A$ be a semisimple MV-algebra. Then a cut $X$ of $A$ is a limit cut if and only if there exists a cut $Y$ of $A$ such that, in $0,1]^{\operatorname{Max}A}$, $\bigvee\widehat{X}=\bigwedge\widehat{Y}^{*}$, where $Y^{*}=\{y^{*}\mid y\in Y\}$. Moreover, $Y$ is a limit cut too.
@@
[14, Corollary 4.6] A semisimple MV-algebra $A$ is lcc if and only if, for all $X,Y\subseteq A$ and $\alpha\in[0,1]^{\operatorname{Max}A}$, $\alpha=\bigvee\widehat{X}=\bigwedge\widehat{Y}$ implies $\alpha\in\widehat{A}$.
@@
  • closed if $f^{\to}(\beta)\in\tau^{*}_{Y}$ for all $\beta\in\tau^{*}_{X}$, and
@@
Let us consider the MV-topological space $(\{x\},[0,1]^{\{x\}})$ on a singleton $\{x\}$. For any $n>1$, all the families $0,1/n]^{\{x\}}$, $([0,1/n]\cap\mathbb{Q})^{\{x\}}$, and $([0,1/n]\setminus\mathbb{Q})^{\{x\}}$ are easily seen to be (non-large) subbases for the given MV-topology. In fact, $([0,1]\setminus\mathbb{Q})^{\{x\}}$ is a base.
@@
[14, Lemma 3.9] If $(X,\tau)$ is an Hausdorff space, then all crisp singletons of $X$ are closed.
@@ 3 MV-lattices and Positive MV-algebras
    $(x\cdot y)\vee(x\wedge y)=x\wedge y$.
@@
Every bounded distributive lattice, with $\vee$ and $\cdot:=\wedge$ , is an MV-lattice.
Given an MV-algebra $(A,\oplus,^{*},0)$ and considering its derived operations, the algebra $(A,\oplus,\odot,\vee,\wedge,0,1)$ is an MV-lattice. Any subalgebra $(B,\oplus,\odot,\vee,\wedge,0,1)$ of such an MV-lattice (i.e., any subset of $A$ containing 0 and $1$ and which is closed under the four binary operations) is called a positive subreduct of the MV-algebra $A$. The following definition is a reformulation of [1, Definition 2.2].
@@
In the case of MV-algebras $\text{\L }_{n+1}$, it was shown in [12] that all of their positive MV-subreducts are exactly the MV-subalgebras, i. e., the $\text{\L }_{k+1}$’s, for all divisor $k$ of $n$.
@@
The set $A=\left\{(0,0),\left(0,\frac{1}{2}\right),(0,1),\left(\frac{1}{2},1\right),(1,1)\right\}\subset\text{\L }_{3}\times\text{\L }_{3}$ is a generating positive MV-algebra of $\text{\L }_{3}\times\text{\L }_{3}$
@@
where, the maximal ideals of $\text{\L }_{3}\times\text{\L }_{3}$ are $I_{1}=\{(0,0),(0,\frac{1}{2}),(0,1)\}$ and $I_{2}=\{(0,0),(\frac{1}{2},0),$ (1,0)}(1,0)\}. Then, $\operatorname{Max}_{\leq}A=\{I_{1}^{A},I_{2}^{A}\}$ where $I_{2}^{A}=\{(0,0)\}\subset I_{1}=I_{1}^{A}$. So, $I_{2}^{A}$ is not a maximal ideal of $A$.
@@
The category of MV-lattices and its morphisms will be denoted by $^{\operatorname{MV}}\!\mathcal{L}$, and we will denote by $\mathcal{MV}_{+}$ and $\mathcal{MV}^{\operatorname{lcc}}_{+}$ its full subcategories of positive MV-algebras and $\operatorname{lcc}$ positive MV-algebras, respectively. Given an MV-lattice $(A,+,\cdot,\vee,\wedge,0,1)$, we consider the set of morphisms from $A$ to $[0,1]$, which is denoted by $H_{A}$ and is equipped with a partial order defined by $f\leq g$ if and only if $f(a)\leq g(a)$ for every $a$ in $A$. For $A$ as in Example 3.7, we have that $H_{A}=\{f_{1},f_{2}\}$ where $f_{i}:A\longrightarrow\text{\L }_{3}\subseteq[0,1]$ is defined by $f_{i}(x_{1},x_{2})=x_{i}$ for $(x_{1},x_{2})\in A$ and $f_{1}\leq f_{2}$. Note that $f_{1}$ and $f_{2}$ are restrictions of homomorphisms from the MV-algebra $\text{\L }_{3}\times\text{\L }_{3}$ to $[0,1]$. This is actually an instance of a general situation, as shown in the following lemma.
@@
The following result provides a necessary and sufficient condition for an MV-lattice $A$ to be embedded in a semisimple MV-algebra. We recall that maximal ideals of any MV-algebra are in bijective correspondence with homomorphisms from the given algebra to $[0,1]$. Therefore, an MV-algebra $B$ is semisimple if and only if the intersection of the kernels of all the homomorphisms from $B$ to $[0,1]$ is the identity congruence $\Delta$.
@@
It is known that an MV-algebra can be generated by various positive algebras. According to Lemma 3.9, the corresponding sets of homomorphisms from these algebras into $[0,1]$ are equipotent. However, these sets are not necessarily isomorphic partially ordered sets, as the following example shows.
@@
Given the MV-algebra $E=\text{\L }_{3}\times\text{\L }_{3}\times\text{\L }_{3}$, we have that the set
@@
    The positive MV-algebra C= B ∪{(0,0, 13), (0,0, 23), (12, 12, 23), (0,12,23), (0,0,1), (0, 12,1), (12, 12,1)} is such that I_1^C=I_1,I_2^C={(0,0,0), (0,0, 13), (0,0, 23), (0,0,1)}, andI_3^C={(0,0,0), (0,12,0), (0,1,0), (12, 12,0), (12,1,0), (1,1,0)}. Moreover $f_{1_{C}}\leq f_{2_{C}}$ but $f_{2_{C}}$ and $f_{3_{C}}$ are not comparable, since f_3(12, 12, 23)=23¿12=f_2(12, 12, 23) and f_3(12, 12,0)=0 ¡ 12=f_2(12, 12,0).
    $H_{C}$$\text{Max}_{\leq}(C)$$f_{2_{C}}$$I_{1}^{C}$$f_{3_{C}}$$I_{3}^{C}$$f_{1_{C}}$$I_{2}^{C}$} So, $H_{A}$ and $H_{C}$ are equipotent sets but are not order isomorphic.
@@
    If $x<y<\frac{1}{2}$, we apply Step 1 and we obtain that $y_{1}\ominus x_{1}\geq 2(y\ominus x)>y\ominus x$ where $y_{1}=ky$ and $x_{1}=kx$. Moreover, since $y_{1}>\frac{1}{2}$, then either $x_{1}\leq\frac{1}{2}$ or $x_{1}>\frac{1}{2}$. If $x_{1}\leq\frac{1}{2}$, by case 1, there is $n_{1}\in\mathbb{N}$ such that $n_{1}x_{1}^{2}=0$ and $n_{1}y_{1}^{2}=1$. That is, there is a term $t(a)=n_{1}(ka)^{2}$ such that $t(x)=0$ and $t(y)=1$. If $x_{1}>\frac{1}{2}$, we apply Step 2 and we obtain y_2 ⊖x_2=y_1^h-x_1^h=h(y_1 ⊖x_1) ≥2(y_1⊖x_1)¿ y_1 ⊖x_1 where $y_{2}=y_{1}^{h}$ and $x_{2}=x_{1}^{h}$ . Note that $x_{2}<\frac{1}{2}$ and again we can apply Step 1 or case 2 according to the possible values of $y_{2}$. Thus we define by recursion a sequence of pairs $(x_{i},y_{i})$ such that for each $i\geq 1$, $y_{i+1}\ominus x_{i+1}\geq 2(y_{i}\ominus x_{i})>y_{i}\ominus x_{i}$. For this reason, for some $n\in\mathbb{N}$, $x_{n}$ and $y_{n}$ will satisfy the conditions of case 1 or 2 and we shall obtain a term $t$ such that $t(x)=0$ and $t(y)=1$.
@@ 4 The extension of Priestley Duality
    $x\nleq y$ implies that there is an increasing clopen set $\alpha$ such that $\alpha(x)=1$ and $\alpha(y)=0$, and
@@
    $(X,\tau)$ has a basis of clopens.
MV-spaces satisfying P2 are called totally order disconnected. It is easy to verify that every totally order disconected MV-space is a Hausdorff MV-space. We will denote by $^{\operatorname{MV}}\!\ \mathcal{T}\!\!\operatorname{op}_{\leq}$ the category of partially ordered MV-spaces whose morphisms are functions that are simultaneously continuous and order-preserving. Moreover, we shall denote by $^{\operatorname{MV}}\!\mathcal{P}\!\!\operatorname{ries}$ the full subcategory of $^{\operatorname{MV}}\!\ \mathcal{T}\!\!\operatorname{op}_{\leq}$ whose objects are Priestley MV-spaces.
@@
Let $(A,+,\cdot,\vee,\wedge,0,1)$ be an MV-lattice and $H_{A}$ be the set of homomorphisms $f:A\longrightarrow[0,1]$. $H_{A}$ has a partial order defined by $f\leq g$ iff $f(a)\leq g(a)$ for all $a\in A$. Now, in order to define an MV-topology on $H_{A}$, let us consider the set $\tilde{A}=\{\tilde{a}\mid a\in A\}$ as in Proposition 3.10, where
@@
The set $\tilde{A}$ generates a semisimple MV-algebra $B\subseteq[0,1]^{H_{A}}$ which, on its turn, is a basis for an MV-topology $\tau_{A}$ on $H_{A}$ (see [14, Section 4]). This partially ordered MV-space $(H_{A},\tau_{A},\leq)$ generated by $B$ is called the MV-space associated to $A$.
@@
${\operatorname{Clop}_{\uparrow}X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varepsilon_{\operatorname{Clop}_{\uparrow}X}}$$\scriptstyle{\operatorname{id}_{\operatorname{Clop}_{\uparrow}X}}$${\operatorname{Clop}_{\uparrow}X}$${\operatorname{Clop}_{\uparrow}\Upsilon\operatorname{Clop}_{\uparrow}X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Clop}_{\uparrow}\eta_{X}}$ ${\Upsilon A\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta_{\Upsilon A}}$$\scriptstyle{\operatorname{id}_{\Upsilon A}}$${\Upsilon\operatorname{Clop}_{\uparrow}\Upsilon A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Upsilon\varepsilon_{A}}$${\Upsilon A}$
@@
For any MV-lattice $(A,+,\cdot,\vee,\wedge,0,1)$, the associated MV-space $(H_{A},\tau,\leq)$ is compact.
@@
Let $\mathcal{MV}^{\operatorname{lcc}}_{+H}$ be the category of $\operatorname{lcc}$ positive MV-algebras that are $H$-complete and $^{\operatorname{MV}}\!\mathcal{P}\!\!\operatorname{ries}^{\operatorname{lcc}}$ be the category of Priestley MV-spaces whose set of increasing clopen sets is an $\operatorname{lcc}$ positive MV-algebra. Then, thanks to Theorem 4.5 and the last two lemmas, the following result readily follows.